Minutes PSB Upgrade WG Meeting 201, 26th of September 2017

Participants: Simon Albright, Thomas William Birtwistle, Jean-Baptiste Bonnamy, Louis de Mallac, James Dilwyn Devine, Gian Piero Di Giovanni, Tobias Dobers, Eve Fortescue-Beck, Greta Guidoboni, Klaus Hanke, Jan Hansen, Iņigo Lamas Garcia, Bettina Mikulec, Antony Newborough, Fernando Pedrosa, Gabor Petrika, Serge Pittet, Sylvie Prodon, Andrea Santamaria Garcia, Jocelyn Tan.

Indico event: https://indico.cern.ch/event/666898/

1. Approval of minutes

The minutes of the LIU-PSB WG #200 were approved.

2. Communications

Report on the LIU Project meeting:

3. Follow-up of open actions

T. Dobers (2017-08-31) Make a list of starting points of the transfer lines to be sent for approval in an EDMS document.

  • Discrepancies were found in the files, which will be discussed in a meeting with C. Hessler, T. Birtwistle, I. Efthymiopoulos. Action extended to 31/10/2017 .

T. Dobers (2017-09-22) Measure the position of the beam stopper during ITS3.

  • The data will be updated in Geode this week. Action closed .
Assigned to Due date Description State Result  
T.Dobers 2017-09-22 Measure the position of the beam stopper during ITS3 edit

G. Petrika (2017-08-31) Prepare a document summarizing the ventilation needs for LIU-PSB (EDMS document 1844566).

  • A draft has been circulated. All EDMS documents need to be sent to G. P. Di Giovanni. Action extended to 10/10/2017 .

G. Petrika (2017-09-26) Verify the consolidation required for building 141.

  • Action extended to 10/10/2017 .

G. Petrika (2017-09-30) Iterate with integration on the installation of the new piping system for cooling.

  • K. Hanke has already contacted the integration team for a meeting. Action extended to 31/10/2017 .

G. Petrika (2017-09-30) release a new version of the EDMS document 1758628, summarizing the cooling needs for LIU-PSB.

  • They are waiting for some comments and the document will be circulated next week. There are no changes for the PSB but there are some for the POPS-B, which should be included in a different document. Action extended to 10/10/2017 .

New action :

Assigned to Due date Description State Result  
G.Petrika 2017-10-10 Circulate an EDMS document regarding the new needs of the POPS-B edit

G. Petrika (2017-09-30) Check the viability of the removal of the obsolete heater units during the YETS during TS2.

  • Action closed .

Assigned to Due date Description State Result  
G.Petrika 2017-09-30 Check the viability of the removal of the obsolete heater units during the YETS during TS2. edit

New action :

Assigned to Due date Description State Result  
G.Petrika 2018-02-06 Specify the steps to be taken during the removal of the heaters during the YETS Presented in LIU-PSB Meeting #208. edit

L. A. Lopez-Hernandez (2017-09-30) Report about the status of the building 245 modifications.

  • Follow-up by K. Hanke.

A. Newborough (2017-09-30) Provide the final price of the recombination dipoles and correctors.

  • Factor 3 more expensive than expected. Action closed .
Assigned to Due date Description State Result  
A.Newborough 2017-09-30 Provide the final price of the recombination dipoles and correctors. edit

J-M. Lacroix (2017-09-30) Complete integration of PSB-PS transfer line. B.Riffaud (2017-09-30) Provide proposal for the support of the equipment in the BTP line.

  • Follow-up by K. Hanke.

M. Calviani / I. Lamas Garcia (2017-09-26) Prepare the ECR for the PSB scraper/absorber describing both the dismantling of the current equipment and the installation of the new one.

  • Action extended to 15/02/2018 .

4. Current Status of the PSB Scraper Study (Iņigo Lamas Garcia)

  • The scraper will be composed of two masks made out of graphite: one fixed and one mobile.
  • Preliminary thermomechanical simulations show that the temperature reached in the scraper among all scenarios is of 46 °C, which is below the maximum temperature allowed for graphite in vacuum (50 °C). More detailed simulations will be done and presented at the LIU-PSB WG meetings.
  • Concerning vacuum, some iterations have been done based on the work presented in the LIU-PSB WG meeting #183.
  • Iterations have also been done concerning integration and design.
  • The scraper design has been validated impedance-wise.
  • Transport will use systems already existing for other PSB elements.

  • B. Mikulec asked if it was possible to know the position of the smaller, mobile mask. I. Lamas Garcia answered that it is possible. There are two positions referenced with switches (in and out) and the movement is executed with stepper motors equipped with a resolver capable of giving the position. It is, therefore, possible to move the mask to any position between the switches. The motors are very precise, RadHard and have a good reproducibility.
  • B. Mikulec asked if the smaller mask can be displaced by a couple of millimetres, I. Lamas Garcia answered that an adjustment system will be in place with micron resolution.
  • G. P. Di Giovanni asked about a high frequency component arising from the removal of the small mask due to the section change. It was previously discussed that the impedance could be improved by adding an extension of the fixed mask, when the mobile one is in OUT position, to reduce the section change and hence the impedance. I. Lamas Garcia answered that the extended mask is still an option, but it has to be checked if additional constraints are introduced by this extended mask.
  • B. Mikulec asked how will the device be aligned in case of ground movements. T. Dobers answered that they could install survey targets on the girder.
  • K. Hanke commented that the scraper timeline should be communicated to J. Coupard, who is in charge of drafting the LS2 installation readiness EDMS document, presented in the LIU-PSB WG meeting #199.
  • B. Mikulec commented that the removal of the beam scope window should be added to the ECR. T. Birtwistle commented that the deadline for the ECR is April 2018. The associated action will be set for February 2018.
  • B. Mikulec suggested to keep the current vacuum chamber as a replacement in case of accident.
  • B. Mikulec commented that the functions specification document for the scrapers (EDMS 1578463) was rewritten by G.P Di Giovanni and is being circulated internally this week. It will be sent for approval after revision from the BE/ABP group.

5. AOB

EVM Update (Sylvie Prodon)

  • The schedule variance has globally gone down from 2.73 to 1.85 months due to re-profiling and reporting.
  • Regarding the CERN Expenditure Tracking for the LIU-PSB project, there is currently 54% of charges and 88% of commitment versus the 55% of charges and 84% commitment objectives for the 1st of October 2017.
  • The only groups low in commitment are EN/HE and SMB/SE due to activities scheduled for the end of the year.
  • The data should be updated by the end of this week for budget estimations for the upcoming years.

Injection Q-strips (Louis de Mallac)

  • The requirements of the power converters are exceeded in terms of voltage and power since the ramping up to 2 GeV was not taken into account in the initial functional specification.
  • Three solutions are presented:
    • Use new power converters with higher power and bandwidth.
    • Modify the circuit layout by splitting the QCF and QCD chains to reduce voltage requirements.
    • Modify the current waveforms by reducing the ramping rates of the MPS, which will reduce the induced voltage in the Q-strips.


  • B. Mikulec commented that the third option would be easy to implement.
  • It was agreed that an MD concerning option 3 will be carried out during the YETS, with the installation of one power converter.
  • B. Mikulec remarked that the final magnetic cycles will not be available for the MD and that only the RMS current at injection can be achieved now.
  • The installation of a new cable will be needed, which needs to be added to the cabling request by Georgi Minchev Georgiev. A slot could be reserved during hardware commissioning for a powering test.
  • If option 3 works it will be implemented. The rest of the power converters need to be ready for installation by the end of 2019 for LS2.
  • A. Newborough asked if an off-the-shelf power converter could be used for the tests, even if over-dimensioned. S. Pittet answered that there are no off-the-shelf power converters with such a bandwidth and power compatible with the space requested for the installation.

6. Tour de table

  • J. Hansen commented that their team does not know which drawing layout to use for the main bending magnets. A. Newborough answered that their designer left one year ago and that the drawings are not necessarily correct.
New action :
Assigned to Due date Description State Result  
K.Hanke 2017-10-31 Attend an integration meeting concerning the layout drawings for the main bending magnets edit

Next LIU-PSB meeting: 10th of October on the BTP line status by B. Riffaud and J.M. Lacroix (not confirmed).

-- AndreaSantamariaGarcia - 2017-09-27

Edit | Attach | Watch | Print version | History: r10 < r9 < r8 < r7 < r6 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r10 - 2018-02-19 - unknown
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSBUpgrade All webs login

This site is powered by the TWiki collaboration platform Powered by PerlWeb page setup by Thomas Hermanns